# Cosmic observation of a model in the horizon of $f(Q,C)$-gravity

### v1

@@ Abstract
We develop a cosmological model in $f(Q,C)$ gravity within the framework of symmetric teleparallel geometry. In addition to the non-metricity scalar $Q$, our formulation includes the boundary term $C$, which accounts for its deviation from the standard Levi-Civita Ricci scalar $R^{*}$ in the Lagrangian. We derive the field equations for the metric and affine connection, employ them within a cosmological setting, and a vanishing affine connection to derive modified Friedmann equations. We use the latest observational dataset OHD in the redshift range $z\in[0,2.36]$, Pantheon + SH0ES in the redshift range $z\in(0.01,2.26)$, BAO, and the joint datasets OHD + Pantheon + SH0ES and OHD + Pantheon + SH0ES + BAO to constrain the model parameters of our model by employing Markov Chain Monte Carlo (MCMC) method to minimize the $\chi^{2}$ term. Using the constrained free model parameters we examine the behaviour of different physical parameters of the model and also, we notice that the model transits from deceleration to acceleration. Finally, we observe that the model demonstrates an expanding quintessence dark energy model and converges to the $\Lambda$CDM in later times.
@@ I Introduction
To address theoretical hitches in general relativity (GR), like cosmological constant problem, and non-renormalizability, and to reconcile its predictions with observational data, numerous theories of gravitational modification have been introduced. There are various approaches to developing these modifications. One approach is to extend the standard curvature formulation of gravity, leading to theories like $f(R)$ gravity, $f(G)$ gravity, Lovelock gravity, and others. Another approach is based on the torsional formulation of gravity, resulting in theories such as $f(T)$ gravity and $f(T,T_{G})$ gravity [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18]. Based on curvature, Einstein’s theory of general relativity (GR) has been remarkably successful due to its strong theoretical coherence and close alignment with experimental observations. However, this success has overshadowed the recognition of two alternative, equally valid formulations of GR in a spacetime where curvature is absent. In such contexts, gravity can be entirely explained by either torsion or the non-metricity of the spacetime.
@@
This occurs because the non-metricity scalar $Q$ and the torsion scalar $T$ differ from the conventional Levi-Civita Ricci scalar $R^{*}$111all quantities indicated with a (*) symbol are computed using the Levi-Civita connection $\Gamma^{*}$ of general relativity by a total divergence term: specifically, $R^{*}=-T+B$ and $R^{*}=Q+C$. As a result, the arbitrary functions $f(R^{*})$, $f(T)$, and $f(Q)$ are no longer equivalent to a total derivative. Furthermore, scalar fields can be incorporated into these frameworks, resulting in scalar-tensor, scalar-torsion, and scalar-non-metricity theories [19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 35, 34, 36, 37, 38, 39, 40, 41, 42, 43, 44].
@@
In the context of teleparallel gravities, the boundary term $B$ can be incorporated into the Lagrangian, leading to $f(T,B)$ theories, which naturally display a more complex phenomenology. However, this work aims to explore this direction by developing $f(Q,C)$ gravity and applying it within a cosmological framework. Within the framework of metric-affine theories, symmetric teleparallel gravity and its variations are gaining prominence in the ongoing efforts to develop a unified theory of gravity. Here, “symmetric” means that parallelograms formed by parallel transport of two vectors do not close, indicating zero torsion. On the other hand, “teleparallel” means that a vector transported around a closed loop remains unchanged, indicating zero curvature. This method describes the geometric and dynamical effects using the non-metricity of the affine connection, characterized by zero curvature and torsion. Unlike in General Relativity, which uses the Levi-Civita connection, the metric tensor and the affine connection in this framework are treated as separate entities that interact through field equations. A major difference between $f(R^{*})$ and $f(Q)$ theories is that $f(Q)$ is a second-order theory like General Relativity (GR), whereas $f(R^{*})$ theory leads to fourth-order field equations. One method to increase the order in symmetric teleparallel theories is by adding higher-order terms such as $\nabla Q$ and $\nabla^{k}Q$ to the Lagrangian. However, a more intuitive approach is to incorporate the boundary term $C$ into the Lagrangian, as shown in studies [1, 4, 45, 37, 34]. It is important to note that this $C$ is not the same as the Gibbons-Hawking-York boundary term, which is required for a well-defined variational formulation of GR in spacetimes with boundaries. The resulting theory is commonly referred to as the symmetric teleparallel equivalent of the $f(R^{*})$ theory because with a specific choice of the function $f(Q,C)=f(Q+C)$, we can derive the $f(R^{*})$ theory [46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56].
@@
The motivation behind this work is the model discussed by Avik De et al. [57]. In that article, the authors discussed possible findings for the future, and in our work, we included both observational and perturbation analyses. Our work is organized in the following manner: In Sec. II, we obtain the field equation with vanishing affine connection $\Gamma^{k}_{ij}=0$ and choose a suitable ansatz of scale factor for further analysis. In the next section III, we study the different observational datasets and methodologies to constrain the free model parameters. In Sec. IV, we explore the physical and dynamical cosmic parameters along with energy conditions, statefinder diagnostic technique, and perturbation analysis. Finally, we conclude our work in the last section V.
@@ II $f(Q,C)$ gravity with FLRW Cosmology
$P^{k}_{ij}=\frac{1}{4}(-2L^{k}_{ij}+Q^{k}g_{ij}-\tilde{Q}^{k}g_{ij}-\delta^{k}_{(i}Q_{j)}).$ (4)
@@
Now, the introduction of the constraints on torsion and curvature i.e. $T^{\lambda}_{ij}:=2\Gamma^{\lambda}_{[ji]}=0$ and $R^{k}_{i\lambda j}:=2\partial_{[\lambda}\Gamma^{k}_{i|j]}+2\Gamma^{k}_{\sigma[\lambda}\Gamma^{\sigma}_{i|j]}=0$, one can derive additional relationships:
@@
Since $Q^{\lambda}-\tilde{Q}^{\lambda}=L^{\lambda}-\tilde{L}^{\lambda}$, following from the earlier statement, the boundary term is also defined as
@@
Here, $f_{C}$ and $f_{Q}$ are the derivatives of $f$ with respect to $C$ and $Q$, respectively. Now, it is crucial to note that unlike in General Relativity, the teleparallel theory features an affine connection separate from the metric tensor, both of which are treated as dynamic variables. Consequently, varying the action concerning the affine connection yields the connection field equation $(\nabla_{i}-\tilde{L}_{i})(\nabla_{j}-\tilde{L}_{j})[4(f_{Q}-f_{C})P^{ij}_{k}]=0$, in the absence of the hyper-momentum tensor.
We explore $f(Q,C)$ gravity within a cosmological context, specifically focusing on $f(Q,C)$ gravity. We investigate a homogeneous and isotropic flat FLRW spacetime, described by the line element.
@@
where $a(t)$ is the scale factor of the Universe. Within the framework of $f(Q,C)$ gravity, we obtain the effective dark-energy sector with energy-momentum tensor is obtained as
@@
The cosmological uniformity and isotropy of the FLRW metric (13) are exemplified by spatial rotations and translations. An asymmetric teleparallel affine connection, which is free from torsion and curvature, exhibits both spherical and translational symmetries. This implies that the derivative operations of the connection coefficients concerning the vector fields generating spatial rotations and translations vanish. Three specific types of affine connections demonstrate these symmetries. We consider the case where the affine connection vanishes, $\Gamma^{k}_{ij}=0$, and when fixing the coincident gauge we obtained the Friedmann-like equations as [43, 59]
@@
where ${}Q=-6H^{2},~{}R^{*}=6(2H^{2}+\dot{H}),~{}C=R^{*}-Q=6(3H^{2}+\dot{H})$.
We observed that the equations (17) and (18) are reliant on the function $f(Q,C)$ and its derivatives. One might consider selecting functions that yield GR equations of motion when additional terms vanish to incorporate GR solutions alongside effects applicable during early or late times. A possible choice is the function $f(Q,C)=Q^{\mu}+\zeta C^{\nu}$, where $\mu$, $\nu$ and $\zeta$ are constants. For simplicity, we set $\mu=2$, $\nu=1$, and $\zeta=2$ and explore the behavior of dynamical and physical parameters. Also, by substituting $\zeta=0$ and $\mu=1$ one can obtain the standard GR equation for matter dominant phase.
Now, for further analysis, we need to solve the field equations. But, since there are three unknowns in the field equations (17) and (18) namely, $H$, $\rho_{\text{DE}}$ and $p_{\text{DE}}$, thus we need one additional equation to examine the behavior the model. In this work, we restrict the scale factor in the shape of the hybrid expansion law. This form triggers a transition in the Universe from a decelerating phase to an accelerating phase, and it is defined as
@@
where $\alpha$ and $\beta$ are the positive constants and $t_{0}$ denotes the present cosmic time. In the literature, it is recognized that this formulation encompasses both power-law cosmology and exponential law cosmology as special cases. When $\alpha=0$, the hybrid expansion law in (19) simplifies to the exponential law and when $\beta=0$, it reduces to the power-law. In different modified theories of gravity, hybrid expansion law has been discussed previously [60, 61].
Now, the Hubble parameter $H$ can be obtained using $H=\frac{\dot{a}}{a}$ as
@@
Here, we constrain the values of $\alpha$ and $\beta$ using the observational datasets and for that purpose, we need to write the value of the Hubble parameter in terms of redshift ($z$). Thus, with the relation $a=\frac{a_{0}}{1+z}$ (at present $a_{0}=1$) and equation (19), one can obtain the value of $t$ in terms of $z$ and then $H$ in terms of $z$ as
@@
where $W$ is the Lambert function, $H_{0}$ is the present value of the Hubble parameter which will also be constrained by observational datasets. In the next section, we use three observational datasets to find the optimal values of the free model parameters $H_{0}$, $\alpha$, and $\beta$. Markov Chain Monte Carlo (MCMC) method is used for this purpose and contours have been plotted using the emcee library in Python language.
@@ III Observational datasets and Methodology
Recent advances in observational cosmology have significantly deepened our comprehension of cosmic evolution, spanning from ancient origins to contemporary developments. These advancements have elucidated the characteristics of dark matter and dark energy, as well as the formation of cosmic structures. Over the past three decades, observational studies in cosmology, driven by instruments such as the Hubble Space Telescope, have generated a vast array of data. Key datasets include the Sloan Digital Sky Survey, which charts the distribution of galaxies and captures current cosmic variations; cosmic microwave background radiation, validating the Big Bang theory; quasars, providing insights into intervening matter between observers and these distant phenomena; and baryon acoustic oscillations, offering measurements of large-scale structures crucial for understanding dark energy. Additionally, supernovae type Ia known as standard candles, is pivotal in determining cosmic distances. This study employs error bar plots of Hubble and SNeIa datasets to compare our model with the $\Lambda$CDM model. Using scipy optimization techniques in Python language, we constrain model parameters and predict the global minimum of the Hubble function as described in Eq. (22). Noticeable fluctuations within the diagonal elements of the covariance matrix related to these parameters are observed. By utilizing these measurements as averages and applying a Gaussian prior with a fixed dispersion value $\sigma=1$, we utilize the emcee module of Python for mathematical research and numerical analysis. This methodology allows us to explore the parameter space surrounding local minima. Results are depicted as two-dimensional contour plots indicating $1\sigma$ and $2\sigma$ uncertainties. I’ve shared the details on the methodologies applied to these datasets below.
@@ III.2 Pantheon + SH0ES Dataset
As technology progresses, our exploration of observational cosmology relies heavily on diverse datasets that illuminate various aspects such as early cosmic evolution, the formation of structures, and the enigmatic properties of the dark Universe, particularly in explaining the accelerated expansion observed in the cosmos using cosmic mechanisms and ray detectors. Among these datasets, the Pantheon + SH0ES sample holds significant prominence, comprising $1701$ data points and serving as a leading collection of observations of supernovae type Ia. The Pantheon + SH0ES dataset spans the spectroscopically observed range of redshifts $z\in(0.01,2.26)$ [76, 77, 78, 79, 80, 81]. The distance moduli are calculated using the equation $\mu_{\text{th},i}=\mu(D_{L})=m-M=5\log_{10}(D_{L})+\mu_{0}$, where $M$ denotes absolute magnitude, $m$ denotes apparent magnitude, and $\mu_{0}$ is a marginalized nuisance parameter. The parameter $\mu_{0}$ can be obtained as $\mu_{0}=5\log_{10}\left(\frac{c}{H_{0}}\text{Mpc}\right)+25$.
@@
where $c$ is the speed of light. The agreement between the Pantheon + SH0ES data and the model predictions is quantified using the $\chi^{2}$ function as
@@
where $\sigma_{\mu}(z_{i})$ denotes the standard error associated with the observed distance modulus $\mu_{\text{obs}}(z_{i})$.

### v2

@@ Abstract
In this work, we developed a cosmological model in $f(Q,C)$ gravity within the framework of symmetric teleparallel geometry. In addition to the non-metricity scalar $Q$, our formulation includes the boundary term $C$, which accounts for its deviation from the standard Levi-Civita Ricci scalar $R^{*}$ in the Lagrangian. We derived the field equations for the metric and affine connection, employed them within a cosmological setting, and a vanishing affine connection to derive modified Friedmann equations. We used the latest observational dataset OHD in the redshift range $z\in[0,2.36]$, Pantheon + SH0ES in the redshift range $z\in(0.01,2.26)$, BAO, and the joint datasets OHD + Pantheon + SH0ES and OHD + Pantheon + SH0ES + BAO to constrain the parameters of our model by employing Markov Chain Monte Carlo (MCMC) method to minimize the $\chi^{2}$ term. Using the constrained free model parameters, we carefully analyzed the behavior of different physical parameters and verified that the model transits from deceleration to acceleration. Finally, we observed that the model demonstrates an expanding quintessence dark energy model and converges to the $\Lambda$CDM in later times.
@@ I Introduction
Numerous theories of gravitational modification have been introduced to address theoretical hitches in general relativity (GR), like cosmological constant problem and non-renormalizability, and to reconcile its predictions with observational data. There are various approaches to developing these modifications. One approach is to extend the standard curvature formulation of gravity, leading to theories like $f(R)$ gravity, $f(G)$ gravity, Lovelock gravity, and others. Another approach is based on the torsional formulation of gravity, resulting in theories such as $f(T)$ gravity and $f(T,T_{G})$ gravity [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18]. Based on curvature, Einstein’s theory of general relativity (GR) has been remarkably successful due to its strong theoretical coherence and close alignment with experimental observations. However, this success has overshadowed the recognition of two alternative, equally valid formulations of GR in a spacetime where curvature is absent. In such contexts, gravity can be entirely explained by either torsion or the non-metricity of the spacetime.
@@
This is a consequence of the non-metricity scalar $Q$ and the torsion scalar $T$ differ from the conventional Levi-Civita Ricci scalar $R^{*}$111all quantities indicated with a (*) symbol are computed using the Levi-Civita connection $\Gamma^{*}$ of general relativity by a total divergence term: specifically, $R^{*}=-T+B$ and $R^{*}=Q+C$. As a result, the arbitrary functions $f(R^{*})$, $f(T)$, and $f(Q)$ are no longer equivalent to a total derivative. Furthermore, scalar fields can be incorporated into these frameworks, resulting in scalar-tensor, scalar-torsion, and scalar-non-metricity theories [19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 35, 34, 36, 37, 38, 39, 40, 41, 42, 43, 44].
@@
In the context of teleparallel gravities, the boundary term $B$ can be incorporated into the Lagrangian, leading to $f(T,B)$ theories, which naturally display a more complex phenomenology. In this work, we intend to explore this direction by developing $f(Q,C)$ gravity and applying it within a cosmological framework. Within the framework of metric-affine theories, symmetric teleparallel gravity and its variations are gaining prominence in the ongoing efforts to develop a unified theory of gravity. Here, “symmetric” means that parallelograms formed by parallel transport of two vectors do not close, indicating zero torsion. On the other hand, “teleparallel” means that a vector transported around a closed loop remains unchanged, indicating zero curvature. This method describes the geometric and dynamical effects using the non-metricity of the affine connection, characterized by zero curvature and torsion. Unlike in General Relativity, which uses the Levi-Civita connection, the metric tensor and the affine connection in this framework are treated as separate entities that interact through field equations. A major difference between $f(R^{*})$ and $f(Q)$ theories is that $f(Q)$ is a second-order theory like General Relativity (GR), whereas $f(R^{*})$ theory leads to fourth-order field equations. One method to increase the order in symmetric teleparallel theories is by adding higher-order terms such as $\nabla Q$ and $\nabla^{k}Q$ to the Lagrangian. However, a more intuitive approach is to incorporate the boundary term $C$ into the Lagrangian, as shown in studies [1, 4, 45, 37, 34]. It is important to note that this $C$ is not the same as the Gibbons-Hawking-York boundary term, which is required for a well-defined variational formulation of GR in spacetimes with boundaries. The resulting theory is commonly referred to as the symmetric teleparallel equivalent of the $f(R^{*})$ theory because with a specific choice of the function $f(Q,C)=f(Q+C)$, we can derive the $f(R^{*})$ theory [46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56].
@@
The motivation behind this work is the model discussed by Avik De et al. [57]. In that article, the authors pointed out some perspectives for further applications of their model, and some of them are addressed here, where we approached both observational and perturbation analyses. Our work is organized in the following nutshell: In Sec. II, we obtain the field equation with vanishing affine connection $\Gamma^{k}_{ij}=0$ and choose a suitable Ansatz of scale factor for further analysis. In the next section III, we study the different observational datasets and methodologies to constrain the free model parameters. In Sec. IV, we explore the physical and dynamical cosmic parameters, energy conditions, statefinder diagnostic technique, and perturbation analysis. Finally, in section V, we present our conclusions and perspectives.
@@ II $f(Q,C)$ gravity with FLRW Cosmology
$P^{k}_{ij}=\frac{1}{4}(-2L^{k}_{ij}+Q^{k}g_{ij}-\tilde{Q}^{k}g_{ij}-\delta^{k}_{(i}Q_{j)})\,,$ (4)
@@
Now, by adding of the constraints on torsion and curvature i.e. $T^{\lambda}_{ij}:=2\Gamma^{\lambda}_{[ji]}=0$ and $R^{k}_{i\lambda j}:=2\partial_{[\lambda}\Gamma^{k}_{i|j]}+2\Gamma^{k}_{\sigma[\lambda}\Gamma^{\sigma}_{i|j]}=0$, one can derive the following equations:
@@
Once $Q^{\lambda}-\tilde{Q}^{\lambda}=L^{\lambda}-\tilde{L}^{\lambda}$, following from the earlier statement, the boundary term is also defined as
@@
Here, $f_{C}$ and $f_{Q}$ are the derivatives of $f$ with respect to $C$ and $Q$, respectively. It is crucial to note that unlike in General Relativity, the teleparallel theory features an affine connection separate from the metric tensor, both of which are treated as dynamic variables. Consequently, varying the action concerning the affine connection yields the connection field equation $(\nabla_{i}-\tilde{L}_{i})(\nabla_{j}-\tilde{L}_{j})[4(f_{Q}-f_{C})P^{ij}_{k}]=0$, in the absence of the hyper-momentum tensor.
We explore $f(Q,C)$ gravity within a cosmological context, specifically focusing on $f(Q,C)$ gravity. We investigate a homogeneous and isotropic flat FLRW spacetime described by the line element.
@@
where $a(t)$ is the scale factor of the Universe. Within the framework of $f(Q,C)$ gravity, we obtain the effective dark-energy sector by considering the energy-momentum tensor as
@@
The cosmological uniformity and isotropy of the FLRW metric (13) are exemplified by spatial rotations and translations. An asymmetric teleparallel affine connection, free from torsion and curvature, exhibits both spherical and translational symmetries. This implies that the derivative operations of the connection coefficients concerning the vector fields generating spatial rotations and translations vanish. Three specific types of affine connections demonstrate these symmetries. We consider the case where the affine connection vanishes, $\Gamma^{k}_{ij}=0$, and when fixing the coincident gauge, we obtained the Friedmann-like equations as [43, 59]
@@
for ${}Q=-6H^{2},~{}R^{*}=6(2H^{2}+\dot{H}),~{}C=R^{*}-Q=6(3H^{2}+\dot{H})$.
We observed that the equations (17) and (18) are reliant on the function $f(Q,C)$ and its derivatives. One might consider selecting functions that yield GR equations of motion when additional terms vanish to incorporate GR solutions alongside effects applicable during early or late times. A possible choice is the function $f(Q,C)=Q^{\mu}+\zeta C^{\nu}$, where $\mu$, $\nu$ and $\zeta$ are constants. For simplicity, we set $\mu=2$, $\nu=1$, and $\zeta=2$ and explore the behavior of dynamical and physical parameters. It is relevant to highlight that such choices generalize this model in respect to general relativity, which can be recovered by setting $\zeta=0$ and $\mu=1$.
Now, for further analysis, we need to solve the field equations. But, once there are three unknowns in equations (17) and (18), namely, $H$, $\rho_{\text{DE}}$ and $p_{\text{DE}}$, we need one additional equation to examine the behavior the model. In this work, we restrict the scale factor in the shape of the hybrid expansion law. This form triggers a transition in the Universe from a decelerating phase to an accelerating phase, and it is defined as
@@
where $\alpha$ and $\beta$ are the positive constants and $t_{0}$ denotes the present cosmic time. In the literature, it is recognized that this formulation encompasses both power-law cosmology and exponential law cosmology as special cases. When $\alpha=0$, the hybrid expansion law in (19) simplifies to the exponential law and when $\beta=0$, it reduces to the power-law [60, 61].
Now, the Hubble parameter $H$ can be obtained by rewriting $H=\frac{\dot{a}}{a}$ as
@@
Here, we constrain the values of $\alpha$ and $\beta$ using the observational datasets and for that purpose, we need to have the Hubble parameter as a function of the redshift ($z$). Thus, with the relation $a=\frac{a_{0}}{1+z}$ (at present $a_{0}=1$) and equation (19), one can obtain the value of $t$ in terms of $z$
@@
where $W$ is the Lambert function, $H_{0}$ is the present value of the Hubble parameter which will also be constrained by observational datasets. In the next section, we use three observational datasets to find the optimal values of the free model parameters $H_{0}$, $\alpha$, and $\beta$. Markov Chain Monte Carlo (MCMC) method is applied for this purpose, and contours were plotted using the emcee library in Python.
@@ III Observational datasets and Methodology
Recent advances in observational cosmology have significantly deepened our comprehension of cosmic evolution, spanning from ancient origins to contemporary developments. These advancements have elucidated the characteristics of dark matter and dark energy, as well as the formation of cosmic structures. Over the past three decades, observational studies in cosmology, driven by instruments such as the Hubble Space Telescope, have generated a vast array of data. Key datasets include the Sloan Digital Sky Survey, which charts the distribution of galaxies and captures current cosmic variations; cosmic microwave background radiation, validating the Big Bang theory; quasars, providing insights into intervening matter between observers and these distant phenomena; and baryon acoustic oscillations, offering measurements of large-scale structures crucial for understanding dark energy. Additionally, supernovae type Ia, known as standard candles, is pivotal in determining cosmic distances. This study employs error bar plots of Hubble and SNeIa datasets to compare our model with the $\Lambda$CDM model. Using optimization techniques in Python, we constrain model parameters and predict the global minimum of the Hubble function as described in Eq. (22). Noticeable fluctuations within the diagonal elements of the covariance matrix related to these parameters are observed. By utilizing these measurements as averages and applying a Gaussian prior with a fixed dispersion value $\sigma=1$, we use the emcee module of Python for mathematical research and numerical analysis. This methodology allows us to explore the parameter space surrounding local minima. Results are depicted as two-dimensional contour plots indicating $1\sigma$ and $2\sigma$ uncertainties. I’ve shared the details on the methodologies applied to these datasets below.
@@ III.2 Pantheon + SH0ES Dataset
As technology progresses, our exploration of observational cosmology relies heavily on diverse datasets that illuminate various aspects, such as early cosmic evolution, the formation of structures, and the enigmatic properties of the dark Universe, particularly in explaining the accelerated expansion observed in the cosmos using cosmic mechanisms and ray detectors. Among these datasets, the Pantheon + SH0ES sample holds significant prominence, comprising $1701$ data points and serving as a leading collection of observations of supernovae type Ia. The Pantheon + SH0ES dataset spans the spectroscopically observed range of redshifts $z\in(0.01,2.26)$ [76, 77, 78, 79, 80, 81]. The distance moduli are calculated using the equation $\mu_{\text{th},i}=\mu(D_{L})=m-M=5\log_{10}(D_{L})+\mu_{0}$, where $M$ denotes absolute magnitude, $m$ denotes apparent magnitude, and $\mu_{0}$ is a marginalized nuisance parameter. The parameter $\mu_{0}$ can be obtained as $\mu_{0}=5\log_{10}\left(\frac{c}{H_{0}}\text{Mpc}\right)+25$.
@@
where $c$ is the speed of light. The agreement between the Pantheon + SH0ES data and the model predictions is quantified by taking the $\chi^{2}$ function as
@@
here $\sigma_{\mu}(z_{i})$ denotes the standard error associated with the observed distance modulus $\mu_{\text{obs}}(z_{i})$.